OpenStudy (anonymous):

Which of the following is equivalent to log9 36? log9/log36 log36/log9 log4 log 1/4

OpenStudy (saifoo.khan):

Do you know about change of base law?

OpenStudy (saifoo.khan):

\[\Large \log_a b = \frac{\log b}{\log a}\]

OpenStudy (anonymous):

Whoo change of base

OpenStudy (anonymous):

Thats what I forget about!

OpenStudy (anonymous):

vaguely..

OpenStudy (anonymous):

Use change of base to get the 2nd one.

OpenStudy (anonymous):

a=9 and b=36